OpenStudy (anonymous):

HELP WITH ALGEBRA 2, PLEASE? Will Fan and give medal. (: 3. The same wax will be used to make each candle, and it costs 3/4 of a penny per cubic centimeter. Write polynomial functions to show the cost of the wax that is used for making each candle for different values of x. I have two different candles. One is called The Brick and has a volume of x^3 - x^2. (^ represents exponent) The other is called The Egyptian and has a volume of 1/3x^2 + 7x.

OpenStudy (anonymous):

@TheSmartOne Can you help?

OpenStudy (anonymous):

@SolomonZelman ?

OpenStudy (solomonzelman):

the volume functions for cancels, are f(x)=x^3-x^2 (Brick) g(x)=\(\large\color{black}{\frac{1}{3}}\)x^2+7x (Egyptian) and the g(x) and f(x) are outputs, of a cubic centimeters of wax per number of candels x, or what?

OpenStudy (anonymous):

I don't know.. I'm still confused by this problem. :(

OpenStudy (anonymous):

@SolomonZelman Could you explain a bit more, please?

OpenStudy (solomonzelman):

you know that \(\large\color{black}{ \rm cm^3~of~wax }\)\(\large\color{black}{ =0.75~c }\) . So, to find the cost of each cancel, I would multiply inside each input times 0.75

OpenStudy (solomonzelman):

(this cost would be in terms of pennies though, don't forget.)

OpenStudy (anonymous):

Alright, so basically all I need to do is multiply 0.75 by each equation? Like, x^3 - x^2 * 0.75 = ?

OpenStudy (solomonzelman):

well, like , \(\large\color{black}{ (x^3 - x^2)\times 0.75 }\)

OpenStudy (anonymous):

Okay. I calculated it by myself as well as on an online calculator. For both answers, I got 3/4 * x^3 -3/4 *x^2. Is that correct? Thank you so much for helping me. (:

OpenStudy (anonymous):

@SolomonZelman